# Knowledge Editing for Large Language Models: A Survey

### v3

@@ 5.4.4. Summary
In this part, we introduce the local modification strategy for pre-trained LLMs for efficient updates of new information without adding new weights or optimizing the whole network. We start by analyzing the pivotal role of the point-wise feedforward layers, i.e., the FFNs, to store the factual information in pre-trained LLMs, with the knowledge neurons associated with the FFN layer thoroughly analyzed. We then discuss the gradient-based and prompt-based strategies that conduct local modification of LLMs, which achieve the modification in a top-down manner based on the output $y$ and in a bottom-up manner with the input prompt $x$, respectively. Nevertheless, the scalability and retainability of local modification methods lack further improvements, as the performance might deteriorate with more edits performed (Meng et al., 2023).
@@ 6. Datasets
Recently, multiple datasets have been established to facilitate the evaluation of KME methods. In this section, we summarize the commonly-used datasets in Table  3 to benefit future KME research. Specifically, these datasets can be divided into two groups: textual-output datasets (i.e., generation tasks) and categorical-output datasets (i.e., classification tasks). The datasets are obtained from a variety of sources, including knowledge graphs, Wikipedia pages, crowd-sourced responses, etc., which are adapted by researchers to fit into the KME setting.
@@ 6.1. Generation Datasets
For generation datasets, the target is in the form of textual content that is required to be generated by LLMs. Serving as pivotal resources to evaluate KME methods, most generation datasets are based on relational knowledge and used for assessing the ability of editing techniques to inject new factual knowledge. This is because relational datasets preserve more definitive answers for each input and thus are more convenient and precise for evaluation. Specifically, these datasets are generally curated from the corresponding relational datasets to encompass diverse relational contexts, ranging from question-answer pairs to intricate multi-hop queries. Therefore, the most prevalent output format is an object to be predicted.
@@
zsRE Relational 244,173 244,173 Factual Statement Object (De Cao et al., 2021; Mitchell et al., 2022a; Meng et al., 2022; Mitchell et al., 2022b; Huang et al., 2023; Hartvigsen et al., 2022; Meng et al., 2023; Lee et al., 2022; Ni et al., 2023)
@@
T-REx-100/-1000 Relational N/A 100/1,000 Factual Statement Object (Dong et al., 2022a; Li et al., 2023c)
CounterFact Relational N/A 21,919 Factual Question Object (Meng et al., 2022; Zheng et al., 2023; Meng et al., 2023; Ni et al., 2023)
@@
MQuAKE-CF/-T Relational N/A 9,218/1,825 Multi-hop Question Object (Zhong et al., 2023)
Hallucination Hallucination N/A 1,392 (Fake) Biography Biography (Hartvigsen et al., 2022)
MMEdit-E-VQA Multimodal 6,346 2,093 Image & Question Answer (Cheng et al., 2023a)
MMEdit-E-IC Multimodal 2,849 1,000 Image Description (Cheng et al., 2023a)
@@
FEVER Fact Checking 104,966 10,444 Fact Description Binary Label (De Cao et al., 2021; Mitchell et al., 2022a; Huang et al., 2023)
@@
SCOTUS Categorization 7,400 931 Court Documents Dispute Topic (Hartvigsen et al., 2022)
  • zsRE (Levy et al., 2017): zsRE is one of the most prevalent Question Answering (QA) datasets extended and adopted by (De Cao et al., 2021; Mitchell et al., 2022a) for KME evaluation. zsRE is suitable for evaluating KME due to its annotations of human-generated question paraphrases, which allow researchers to assess the model resilience to semantically equivalent inputs. In zsRE, each relation is associated with a set of crowd-sourced template questions, such as "What is Albert Einstein’s alma mater?". Each data entry cites a Wikipedia sentence that serves as the factual basis or provenance. The dataset also contains negative examples, which are generated by pairing a valid question with a random sentence.
@@
  • Hallucination (Hartvigsen et al., 2022): Firstly processed in GRACE (Hartvigsen et al., 2022), Hallucination is created from the dataset released in SelfCheckGPT (Manakul et al., 2023), where the authors prompt GPT-3 to generate biographies based on concepts extracted from WikiBio. The sentences are annotated regarding the factual accuracy, and hallucinations in them are identified. Then in GRACE, the authors process this dataset by further extracting Wikipedia summaries from WikiBio and thereby acquire the correct entry of each sentence. In this manner, every edit consists of a potentially false biography generated by GPT-3 as prompt, and a ground truth output, which is the correct next sentence extracted from Wikipedia. There exist 1,392 potential edits for test.
@@
  • MMEdit (Cheng et al., 2023a): This dataset is the first to explore the possibility of editing multimodal LLMs. Specifically, MMEdit consists of two prevalent multimodal tasks: Visual Question Answering (VQA) (Antol et al., 2015) and Image Captioning (Herdade et al., 2019). VQA involves developing algorithms that can analyze an image’s visual content, comprehend questions asked in natural language about the image, and accurately respond to those questions. Image Captioning aims to understand an image and then generate a detailed and coherent natural language description of that image. To create dataset MMEdit, the authors utilize BLIP-2 OPT (Li et al., 2023a) and extract edit data from the evaluation datasets VQAv2 (Goyal et al., 2017) and COCO Caption (Chen et al., 2015), specifically focusing on their suboptimal entries.
@@
  • ECBD (Onoe et al., 2023): Based on the original dataset ECBD (Entity Cloze By Date) (Onoe et al., 2022), the authors process this dataset for a novel task, namely Entity Knowledge Propagation (EKP). The task aimed at updating model parameters to incorporate knowledge about newly emerged entities that are not present in the pre-training data of the language models. For instance, BERT (Devlin et al., 2018), trained in 2018, does not recognize “COVID-19” as it is a more recent entity. The processed dataset aims to provide evaluation for such a task with the help of definition sentences as input to update knowledge about new entities. The entities are taken from date between 2020/01 and 2021/09 to ensure that they are not in training data. Each edit consists of a new entity, a description sentence, a probe sentence, and a ground truth completion
@@ 6.2. Classification Datasets
Classification datasets are also widely adopted to evaluate the effectiveness of KME. These datasets consist of prompt-target pairs, where the target is a discrete label instead of a textual sentence. In the context of KME, these labels help ascertain the alignment of model performance with desired edits, whether they are sentiment-based, fact-checking, or entaiKMEnt-related. The advantages of classification datasets also involve their preciseness in evaluation without the need to define the specific output space. In this section, we summarize notable classification datasets that have been tailored and leveraged for assessing KME techniques as follows:
@@
  • ConvSent (Mitchell et al., 2022b): Firstly processed in SERAC (Mitchell et al., 2022b), ConvSent is used to evaluate the capability of an editor to modify a dialog agent’s sentiment about a particular topic without influencing its responses to other topics. ConvSent is obtained from a list of 15,000 non-numeric entities from zsRE (Levy et al., 2017; De Cao et al., 2021), combined with 989 noun phrases from GPT-3 (Brown et al., 2020) for 15,989 topics. Particularly, for each entity, there are ten positive and ten negative sentiment completions, which can be noisy, from the BlenderBot model with 3B parameters (Roller et al., 2021). The refined sentiment labels are achieved by a sentiment classifier (Heitmann, 2020) pre-trained on RoBERTa (Liu et al., 2019).
  • Bias in Bios (De-Arteaga et al., 2019): Bias in Bios is a dataset originally proposed for fairness-related machine learning, containing approximately 397k short professional biographies of online individuals, which are not relatively famous. Each biographical sentence is assigned an associated occupation label for the described person. To adopt this dataset for evaluating the performance of KME methods, the authors of REMEDI (Hernandez et al., 2023) extract a single sentence, modify it to display only the person’s first name, and then query the language model with the prompt that follows the structure: "Person has the occupation of…". Then they evaluate the relative probabilities of the language model assigned to 28 potential occupations, where the language model is considered to be correct if the ground-truth occupation is ranked top-1.
@@
  • SCOTUS (Hartvigsen et al., 2022): Firstly proposed in GRACE (Hartvigsen et al., 2022), SCOTUS is processed with label shift based on the dataset with the same name from Fairlex (Chalkidis et al., 2022). This classification task is to categorize U.S. Supreme Court documents from various decades into one of 11 topics. The topics are clustered based on the specific matter of dispute, such as Criminal Procedure, Civil Rights, and First Amendment. Due to the evolution of categorization rules over time, the label distributions in this dataset also shift. Specifically, 7.4k cases from 1946-1982 are used for training, and 931 cases from the 1991-2009 period are for test.
@@ 7.1. Question Answering
7.1.1. Background.
Question Answering (QA) is a core NLP task that aims to comprehend queries posed by users in natural language and provide answers based on the encoded knowledge in the pre-trained language model (Shin et al., 2020). Traditional models for QA are generally fixed in their knowledge, capturing only the information available at the training time of (Petroni et al., 2019; Jiang et al., 2020). However, in our dynamic world, new information is generated incessantly, which necessitates the constant update of QA models (Talmor et al., 2018). Fortunately, KME methods enable the modification of QA models to cater to specific questions without disrupting responses to other unrelated inputs. Therefore, with KME strategies, the QA model can be efficiently updated on the run, where the currentness of the model can be guaranteed. Consequently, language model editing techniques have found broad applications across a myriad of QA contexts with potentially distinct requirements (Lee et al., 2022).
7.1.2. Existing Works.
The QA task encompasses various aspects, such as conversational QA, definition-based QA, and notably, relation-based QA (Pandya and Bhatt, 2021). Relation-based QA is primarily adopted as an evaluation benchmark as it necessitates the retrieval of precise real-world facts in response to queries. This particular emphasis on specific information retrieval renders relation-based QA especially conducive to the benefits of KME techniques. For example, PPA (Lee et al., 2022) introduces an innovative task of CuQA (Continuously-updated QA), which intentionally emphasizes recurrent, substantial edits for language models to constantly update them with new information. An important aspect of the CuQA task is to ensure that the existing pre-trained knowledge remains unaltered with the integration of new knowledge. Therefore, this property is one important evaluation to assess model editing in CuQA tasks. In MQuAKE (Zhong et al., 2023), the authors innovatively propose a multi-hop QA task that involves answering questions generated by chains of facts in plain text. Specifically, the task requires edited models to infer implicit relations that can be several hops away from the objects in the edit. For example, when a language model is modified regarding the president of the USA, an ideal model should also authentically alter answers to “Who is the son of the president of the USA”, which is a two-hop relation. Such a task is significantly more challenging as it necessitates the model to alter its reasoning results in addition to the original edit. Nevertheless, the proposed method MeLLo in MQuAKE still exhibits outstanding performance on this difficult task, demonstrating the potential of KME in generalizing edited knowledge to multi-hop relations.
@@ 7.2. Fact Checking
7.2.1. Background.
Fact-checking (FC) is a pivotal task in journalism, information verification, and combating misinformation that aims to scrutinize and affirm the authenticity of claims, statements, or information in news articles, social media, and other media content (Schuster et al., 2021; Galitsky, 2023). In a world overwhelmed with ever-emerging information, fact-checking facilitates the trustworthiness in the sharing of distributed information, promotes information transparency, and aids individuals in making well-informed decisions (Thorne et al., 2018). However, it is crucial to constantly update fact-checking models. For instance, during the COVID-19 pandemic, initial understandings and guidelines about the virus evolved as researchers gathered more data (Shahi et al., 2021). A fact-checking model that cannot adapt to these rapidly changing facts would quickly become outdated and potentially spread misinformation, thereby requiring the application of language model editing. By integrating KME techniques into fact-checking models to consistently update them with the latest information and facts, it becomes possible to ensure the currentness, trustworthiness, and accuracy of the model despite the persistent evolution of information.
7.2.2. Existing Works.
Recently, several works have proposed to apply KME techniques in fact-checking models. In (Zhu et al., 2020), the authors first explore the potential of modifying specific factual knowledge within the transformer backbone of the fact-checking model while ensuring that overall model performance remains intact on facts irrelevant to the editing purpose. Particularly, they identify the critical components within the transformer backbones conducive to effective knowledge modifications. In SERAC (Mitchell et al., 2022b), the authors propose to use evidence gathered from Wikipedia as edit descriptors to update potentially outdated knowledge in the model. The proposed method exhibits significant performance improvements over baselines and can be generalized to other in-scope inputs collected from the same Wikipedia page.

### v4

@@ 5.4.4. Summary
In this part, we introduce the local modification strategy for pre-trained LLMs for efficient updates of new information without adding new weights or optimizing the whole network. We start by analyzing the pivotal role of the point-wise feedforward layers, i.e., the FFNs, to store the factual information in pre-trained LLMs, with the knowledge neurons associated with the FFN layer thoroughly analyzed. We then discuss the groundtruth-based strategies, which achieve the modification in a top-down manner, generally based on least squares objectives computed from the output $y$. We further discuss the prompt-based strategies, which conduct modifications in a bottom-up manner based on the input prompt $x$. Nevertheless, the scalability and retainability of local modification methods lack further improvements, as the performance might deteriorate with more edits performed (Meng et al., 2023).
@@ 6. Datasets
Recently, multiple datasets have been established to facilitate the evaluation of KME methods, and we summarize the commonly-used datasets in Table  2 to benefit future KME research. Specifically, these datasets can be divided into two groups: generation datasets (i.e., textual output) and classification datasets (i.e., categorical output). The datasets are obtained from a variety of sources, including knowledge graphs, Wikipedia pages, crowd-sourced responses, etc., which are adapted by researchers to fit into the KME setting.
@@ 6.1. Generation Datasets
For generation datasets, the target is in the form of textual content that is required to be generated by LLMs. Serving as pivotal resources to evaluate KME methods, most generation datasets are based on relational knowledge and used for assessing the ability of editing techniques to inject new factual knowledge. This is because relational datasets preserve more definitive answers for each input and thus are more convenient and precise for evaluation (Zhang et al., 2024; Yao et al., 2023). Specifically, these datasets are generally curated from the corresponding relational datasets to encompass diverse relational contexts, ranging from question-answer pairs to intricate multi-hop queries. Therefore, the most prevalent output format is an object to be predicted.
@@
zsRE Relational 244,173 244,173 Factual Statement Object (De Cao et al., 2021; Mitchell et al., 2022a; Meng et al., 2022; Mitchell et al., 2022b; Huang et al., 2023; Hartvigsen et al., 2023; Meng et al., 2023; Lee et al., 2022; Ni et al., 2023; Li et al., 2024a; Wang et al., 2024c; Song et al., 2024; Wang et al., 2024b; Gangadhar and Stratos, 2024; Jiang et al., 2024; Yu et al., 2024; Gu et al., 2024; Gupta et al., 2024)
CounterFact Relational N/A 21,919 Factual Question Object (Meng et al., 2022; Zheng et al., 2023; Meng et al., 2023; Ni et al., 2023; Li et al., 2024a; Song et al., 2024; Wang et al., 2024b; Chen et al., 2024; Gangadhar and Stratos, 2024; Sharma et al., 2024; Hu et al., 2024; Gupta et al., 2024; Yoon et al., 2024)
@@
T-REx-100/-1000 Relational N/A 100/1,000 Factual Statement Object (Dong et al., 2022a; Li et al., 2023b)
@@
MQuAKE-CF/-T Relational N/A 9,218/1,825 Multi-hop Question Object (Zhong et al., 2023; Gu et al., 2023; Shi et al., 2024; Wang et al., 2024a; Li et al., 2024b; Jiang et al., 2024)
Hallucination Hallucination N/A 1,392 (Fake) Biography Biography (Hartvigsen et al., 2023; Wang et al., 2024c; Yu et al., 2024)
MMEdit-E-VQA Multimodal 6,346 2,093 Image & Question Answer (Cheng et al., 2023)
MMEdit-E-IC Multimodal 2,849 1,000 Image Description (Cheng et al., 2023)
@@
Conflict Edit Relational N/A 7,500 Factual Statement Object (Li et al., 2024d)
Round Edit Relational N/A 5,000 Factual Statement Object (Li et al., 2024d)
UKE Relational N/A 2,478 Factual Question Object (Wu et al., 2024)
RippleEdits Relational N/A 5,000 Factual Question Object (Cohen et al., 2024; Jiang et al., 2024)
VLKEB Multimodal 5,000 3,174 Image Description (Huang et al., 2024)
MLaKE Multilingual N/A 9,432 Question Answer (Wei et al., 2024)
FEVER Fact Checking 104,966 10,444 Fact Description Binary Label (De Cao et al., 2021; Mitchell et al., 2022a; Huang et al., 2023; Chen et al., 2024)
@@
SCOTUS Categorization 7,400 931 Court Documents Dispute Topic (Hartvigsen et al., 2023; Yu et al., 2024)
  • zsRE (Levy et al., 2017): zsRE is one of the most prevalent Question Answering (QA) datasets extended and adopted by (De Cao et al., 2021; Mitchell et al., 2022a) for KME evaluation. zsRE is suitable for evaluating KME due to its annotations of human-generated question paraphrases, which allow researchers to assess the model resilience to semantically equivalent inputs. In zsRE, each relation is associated with a set of crowd-sourced template questions, such as “What is Albert Einstein’s alma mater?”. Each entry cites a Wikipedia sentence, serving as the factual basis or provenance. The dataset also contains negative examples that are generated by pairing a valid question with a random sentence.
@@
  • Hallucination (Hartvigsen et al., 2023): Firstly processed in GRACE (Hartvigsen et al., 2023), Hallucination is created from the dataset released in SelfCheckGPT (Manakul et al., 2023), where the authors prompt GPT-3 to generate biographies based on concepts extracted from WikiBio. The sentences are annotated regarding the factual accuracy, and hallucinations in them are identified. Then in GRACE, the authors process this dataset by further extracting Wikipedia summaries from WikiBio and thereby acquire the correct entry of each sentence. In this manner, every edit consists of a potentially false biography generated by GPT-3 as the prompt, and a ground truth output, which is the correct next sentence extracted from Wikipedia. There exist 1,392 potential edits for test.
@@
  • MMEdit (Cheng et al., 2023): This dataset is the first to explore the possibility of editing multimodal LLMs. Specifically, MMEdit consists of two prevalent multimodal tasks: Visual Question Answering (VQA) (Antol et al., 2015) and Image Captioning (Herdade et al., 2019). VQA involves developing algorithms that can analyze an image’s visual content, comprehend questions asked in natural language about the image, and accurately respond to those questions. Image Captioning aims to understand an image and then generate a detailed and coherent natural language description of that image. To create dataset MMEdit, the authors utilize BLIP-2 OPT (Li et al., 2023a) and extract edit data from the evaluation datasets VQAv2 (Goyal et al., 2017) and COCO Caption (Chen et al., 2015), specifically focusing on their suboptimal entries.
@@
  • ECBD (Onoe et al., 2023): Based on the original dataset ECBD (Entity Cloze By Date) (Onoe et al., 2022), the authors process this dataset for a novel task, namely Entity Knowledge Propagation (EKP). The task aimed at updating model parameters to incorporate knowledge about newly emerged entities that are not present in the pre-training data of the language models. For instance, BERT (Devlin et al., 2018), trained in 2018, does not recognize “COVID-19” as it is a more recent entity. The processed dataset aims to provide evaluation for such a task with the help of definition sentences as input to update knowledge about new entities. The entities are taken from the date between 2020/01 and 2021/09 to ensure that they are not in training data. Each edit consists of a new entity, a description sentence, a probe sentence, and a ground truth completion.
  • VLKEB (Huang et al., 2024): VLKEB (Large Vision-Language Model Knowledge Editing Benchmark) aims to address the unique challenges of editing large vision-language models, which face additional difficulties due to different data modalities and complex model components with limited data for LVLM editing. VLKEB collects data from the multi-modal knowledge graph MMKG (Liu et al., 2019a) and extends the Portability metric for evaluation. With MMKG, VLKEB binds image data with knowledge entities, which can be used to extract entity-related knowledge for editing data.
  • MLaKE (Wei et al., 2024): MLaKE (Multilingual Language Knowledge Editing) is proposed to evaluate the capability of KME methods in multilingual contexts and multi-hop reasoning across five languages: English, Chinese, Japanese, French, and German. MLaKE aggregates fact chains from Wikipedia in multiple languages and utilizes LLMs to generate questions in both free-form and multiple-choice formats. Notably, existing methods show relatively high generalization for languages within the same language family compared to those from different families. These findings underscore the need for advancements in multilingual knowledge editing.
  • UKE (Wu et al., 2024): UKE (Unstructured Knowledge Editing) is proposed to evaluate the capability of KME methods in updating knowledge based on unstructured texts. Updating LLMs with texts appears to be a more realistic application, which is also more complex and difficult. The authors leverage subjects and objects in Wikidata (Vrandečić and Krötzsch, 2014) and retrieve the corresponding Wikipedia article summaries as unstructured texts. The authors also utilize LLMs to generate summaries for edits in two existing datasets, CounferFact (Meng et al., 2022) and MQuAKE-CF (Zhong et al., 2023), to obtain unstructured texts.
  • RippleEdits (Cohen et al., 2024): This dataset proposes a novel evaluation criterion, which assesses the performance of KME methods on additional edits brought by an existing edit. In particular, injecting new knowledge (e.g., “Jack Depp is the son of Johnny Depp”) introduces a “ripple effect,” which necessitates the model to update related knowledge as well (e.g., “Jack Depp is the sibling of Lily-Rose Depp”). Based on this, the authors construct RippleEdits, consisting of 5,000 edits with various types of ripple effects.
  • Conflict/Round Edit (Li et al., 2024d): This dataset pioneers in investigating the potential side effects of KME methods for LLMs. The proposed dataset and evaluation metrics underline two primary concerns: (1) Knowledge Conflict: Modifying sets of logically conflicting facts can amplify the existing inconsistencies within LLMs. (2) Knowledge Distortion: Altering model parameters to update factual knowledge can permanently disrupt the inherent knowledge framework of LLMs. The dataset is constructed from WikiData (Vrandečić and Krötzsch, 2014) with specific logical rules.
@@ 6.2. Classification Datasets
Classification datasets are also widely adopted to evaluate the effectiveness of KME. These datasets consist of prompt-target pairs, where the target is a discrete label instead of a textual sentence. In the context of KME, these labels help ascertain the alignment of model performance with desired edits. The advantages of classification datasets also involve their preciseness in evaluation without the need to define the specific output space. In this section, we summarize notable classification datasets that have been tailored and leveraged for assessing KME techniques as follows:
@@
  • ConvSent (Mitchell et al., 2022b): Firstly processed in SERAC (Mitchell et al., 2022b), ConvSent is used to evaluate the capability of an editor to modify a dialog agent’s sentiment about a particular topic without influencing its responses to other topics. ConvSent is obtained from a list of 15,000 non-numeric entities from zsRE (Levy et al., 2017; De Cao et al., 2021), combined with 989 noun phrases from GPT-3 (Brown et al., 2020) for 15,989 topics. Particularly, for each entity, there are ten positive and ten negative sentiment completions, which can be noisy, from the BlenderBot model with 3B parameters (Roller et al., 2021). The refined sentiment labels are achieved by a sentiment classifier (Heitmann, 2020) pre-trained on RoBERTa (Liu et al., 2019b).
  • Bias in Bios (De-Arteaga et al., 2019): Bias in Bios is a dataset originally proposed for fairness-related machine learning, containing approximately 397k short professional biographies of online individuals, which are not relatively famous. Each biographical sentence is assigned an associated occupation label for the described person. To adopt this dataset for evaluating the performance of KME methods, the authors of REMEDI (Hernandez et al., 2023) extract a single sentence, modify it to display only the person’s first name, and then query the language model with the prompt that follows the structure: “Person has the occupation of…”. Then they evaluate the relative probabilities of the language model assigned to 28 potential occupations, where the language model is considered to be correct if the ground-truth occupation is ranked top-1.
@@
  • SCOTUS (Hartvigsen et al., 2023): Firstly proposed in GRACE (Hartvigsen et al., 2023), SCOTUS is processed with label shift based on the dataset with the same name from Fairlex (Chalkidis et al., 2022). This classification task is to categorize U.S. Supreme Court documents from various decades into one of 11 topics. The topics are clustered based on the specific matter of dispute, such as Criminal Procedure, Civil Rights, and First Amendment. Due to the evolution of categorization rules over time, the label distributions in this dataset also shift. Specifically, 7.4k cases from 1946-1982 are used for training, and 931 cases from the 1991-2009 period are for test.
@@ 7.1. Question Answering
Background. Question Answering (QA) is a core NLP task that aims to comprehend queries posed by users in natural language and provide answers based on the encoded knowledge in the pre-trained language model (Shin et al., 2020). Traditional models for QA are generally fixed in their knowledge, capturing only the information available at the training time of (Petroni et al., 2019; Jiang et al., 2020). However, in our dynamic world, new information is generated incessantly, which necessitates the constant update of QA models (Talmor et al., 2018). Fortunately, KME methods enable the modification of QA models to cater to specific questions without disrupting responses to other unrelated inputs. Therefore, with KME strategies, the QA model can be efficiently updated on the run, where the currentness of the model can be guaranteed. Consequently, language model editing techniques have found broad applications across a myriad of QA contexts with potentially distinct requirements (Lee et al., 2022).
Existing Works. The QA task encompasses various aspects, such as conversational QA, definition-based QA, and notably, relation-based QA (Pandya and Bhatt, 2021). Relation-based QA is primarily adopted as an evaluation benchmark as it necessitates the retrieval of precise real-world facts in response to queries. This particular emphasis on specific information retrieval renders relation-based QA especially conducive to the benefits of KME techniques. For example, PPA (Lee et al., 2022) introduces an innovative task of CuQA (Continuously-updated QA), which intentionally emphasizes recurrent, substantial edits for language models to constantly update them with new information. An important aspect of the CuQA task is to ensure that the existing pre-trained knowledge remains unaltered with the integration of new knowledge. Therefore, this property is one important evaluation to assess model editing in CuQA tasks. In MQuAKE (Zhong et al., 2023), the authors innovatively propose a multi-hop QA task that involves answering questions generated by chains of facts in plain text. Specifically, the task requires edited models to infer implicit relations that can be several hops away from the objects in the edit. For example, when a language model is modified regarding the president of the USA, an ideal model should also authentically alter answers to “Who is the son of the president of the USA”, which is a two-hop relation. Such a task is significantly more challenging as it necessitates the model to alter its reasoning results in addition to the original edit. Nevertheless, the proposed method MeLLo in MQuAKE still exhibits outstanding performance on this difficult task, demonstrating the potential of KME in generalizing edited knowledge to multi-hop relations.
@@ 7.2. Fact Checking
Background. Fact-checking (FC) is a pivotal task in journalism, information verification, and combating misinformation that aims to scrutinize and affirm the authenticity of claims, statements, or information in news articles, social media, and other media content (Schuster et al., 2021; Galitsky, 2023). In a world overwhelmed with ever-emerging information, fact-checking facilitates the trustworthiness in the sharing of distributed information, promotes information transparency, and aids individuals in making well-informed decisions (Thorne et al., 2018). However, it is crucial to constantly update fact-checking models. For instance, during the COVID-19 pandemic, initial understandings and guidelines about the virus evolved as researchers gathered more data (Shahi et al., 2021). A fact-checking model that cannot adapt to these rapidly changing facts would quickly become outdated and potentially spread misinformation, thereby requiring the application of language model editing. By integrating KME techniques into fact-checking models to consistently update them with the latest information and facts, it becomes possible to ensure the currentness, trustworthiness, and accuracy of the model despite the persistent evolution of information.
Existing Works. Recently, several works have proposed to apply KME techniques in fact-checking models. In (Zhu et al., 2020), the authors first explore the potential of modifying specific factual knowledge within the transformer backbone of the fact-checking model while ensuring that overall model performance remains intact on facts irrelevant to the editing purpose. Particularly, they identify the critical components within the transformer backbones conducive to effective knowledge modifications. In SERAC (Mitchell et al., 2022b), the authors propose to use evidence gathered from Wikipedia as edit descriptors to update potentially outdated knowledge in the model. The proposed method exhibits significant performance improvements over baselines and can be generalized to other in-scope inputs collected from the same Wikipedia page.